OpenStudy (anonymous):

How come that the answer in this question became x=6 and x= -4 ?? Plss. Answer me f(x)= x^2 + 3x + 4 , g(x)= 5x + 28 x^2+3x+4=5x+28 x^2-2x-24=0 (x-6)(x+4)=0 x=6, x=-4 Pls. Answer me :DDD

OpenStudy (anonymous):

f(x)=g(x) will give you first equality

OpenStudy (anonymous):

next - the solution of it... what else to explain?

OpenStudy (anonymous):

because x-6=0 yields x=6 and x+4=0 yields x=-4 after you factor set each factor equal to 0 and solve

OpenStudy (anonymous):

If you place those 6 or -4 in for x at any point in your equation both side should equal zero provided you did it correctly

OpenStudy (anonymous):

Thanks :))))))))